# Federated Automated Feature Engineering

### v2

@@ 2 Introduction
Federated learning (FL) is a privacy-preserving machine learning technique that has gained traction in recent years. The main requirement in federated learning is that data belonging to client devices cannot be directly sent to the server to build a global model. Instead, the typical approach is for each client to build a local model with only their own local data, then the clients can send their model weights to the server which aggregates these weights in various ways. Many of the key challenges in federated learning revolve around minimizing the number of communications between the clients and server and maintaining good performance when the data across clients are non-IID [fl].
There are three federated learning settings that correspond to how the data is gathered across clients. In horizontal federated learning, each client has a subset of samples and each of these samples has all of the features [fl]. In vertical federated learning, each client has the full set of samples but only a subset of the features [fedbcd]. Finally, in hybrid federated learning, each client only has a subset of samples and each of these samples only has a subset of features [hyfdca][hyfem][hybrid-graph]. Each of these settings requires different algorithmic approaches.
This paper focuses on automated feature engineering in the context of federated learning, where clients are not able to send data to the server, but we still build informative new features from the original raw features. We introduce algorithms for all three FL settings: horizontal, vertical, and hybrid. The horizontal algorithm we propose extends IIFE [iife] to the federated setting, which we term Fed-IIFE. The hybrid and vertical FL methods hinge on homomorphic encryption and differential privacy to safely form combinations of features that reside on different clients, along with a feature selection scheme to determine which features to keep. All of the algorithms are iterative in that they gradually construct more and more complicated engineered features each iteration.
@@ 3 Related Work
There is significant work existing on automated feature engineering. The main focus of most algorithms is on cleverly navigating the massive search space of possible combinations of features for high order engineered features. OpenFE uses a very fast boosting technique based on gradient boosted trees to quickly evaluate all possible feature combinations [openfe]. AutoFeat enumerates all possible combinations up to a specified order, then uses a fast feature selection technique to select the best performing features [autofeat]. IIFE tackles this problem by using interaction information to determine the best synergizing feature pairs and only exploring the best pairs of features, then iteratively builds higher order features [iife]. EAAFE uses a genetic algorithm approach to search the best engineered features [eaafe]. DIFER uses a deep learning approach to find representations of the engineered feature strings and their predicted validation scores [difer]. Furthermore, very recently there have been a focus on using Large Language Models to build engineered features in an informed fashion based on feature descriptions [caafe].
@@
Federated learning also has extensive algorithmic work in addition to interesting theoretical results. Federated averaging is the original FL algorithm that is designed for the horizontal setting [fedavg]. Further improvements for heterogeneous data exist, such as SCAFFOLD [scaffold] and RadFed [radfed]. The vertical FL setting has also had significant work in recent years [fedbcd]. The hybrid federated learning case where clients have subsets of both samples and features has far fewer algorithms, but has recently had more work exploring this case [hyfdca][hyfem][hybrid-graph].
@@
Despite the great deal of work in federated learning and automated feature engineering, there is relatively limited work on automated feature engineering in the federated learning setting. All of the existing work at the time of this writing focuses on the vertical case, where each client has different features, and all clients have the same samples. There exists work in AutoML in a federated setting, however automated feature engineering is not one of the aspects discussed in this work [autofl]. FLFE is an algorithm for vertical federated AutoFE that uses quantile sketching to encode features and uses deep neural networks to predict if a particular transformation of features is useful [vertfefl]. The neural networks are trained on a number of datasets, but these datasets are not shared, so it is impossible to reproduce their results. Since these datasets are actually used to build the feature engineering framework and not just used for evaluating the algorithm, it is necessary to have these datasets to properly evaluate the algorithm. Finally, homomorphic encryption is used when constructing the best candidate features for building the final model. However, this work shows only small gains compared to using only the raw features, does not compare to AutoFE methods in the centralized setting, and does not provide any source code making reproducibility very difficult. Furthermore, clients may be able to reconstruct the features of other clients from the decrypted engineered features; we address this in our work by utilizing differential privacy. We believe this makes the work very difficult to reproduce and compare. It is clear that there is a large research gap in all three settings of federated AutoFE: horizontal, vertical, and hybrid. Our work provides algorithms for all of these settings.
@@ 4 Algorithms
Fortunately, extending IIFE to the FL setting is relatively straightforward and results in good empirical performance compared to centralized IIFE, which is one of the top-performing AutoFE algorithms at the time of this writing [iife]. We detail this algorithm, Fed-IIFE, in Section 4.1. The cases of vertical and hybrid FL are more complicated than horizontal FL and are discussed in Section 4.2.
@@ 4.1 Horizontal Setting
We extend IIFE to the FL setting by replacing some key aspects of IIFE with federated versions. We call this new FL version Fed-IIFE. Refer to the full algorithmic description of IIFE in [iife]. We replace computations of interaction information with federated interaction information described in Algorithm 1. The key change is that interaction information is computed separately on each client. The clients send these interaction information scores to the server; the server then averages the contributions from all of the clients. The averaging can either be the simple average or a weighted average where the weights are the proportion of samples belonging to the client. We use the weighted average in our experiments. The other key change from IIFE to Fed-IIFE is that all model evaluations are performed in a federated fashion. Candidate transformations are computed on each client separately and then evaluated across all clients using any horizontal FL algorithm. In our experiments, we use FedAvg [fedavg] for horizontal FL training.
@@ 4.2 Vertical and Hybrid Settings
Vertical AutoFE algorithms must significantly differ from centralized and horizontal FL settings due to the fact that each client holds different feature spaces. The key aspect of AutoFE is combining different features together in ways that improve downstream model performance. However, directly sharing features on different clients with each other or a central server directly violates the requirements of FL. Furthermore, many vertical and hybrid FL algorithms require homomorphic encryption to protect privacy leakages when sharing intermediate information (such as inner products) for computing gradients [hyfdca]. This is computationally expensive which severely limits the number of model training/evaluations that can be performed (AutoFE algorithms typically require a very large number of such trainings/evaluations). In fact, we show in Section 5.2.1 that homomorphic encryption is the main computational bottleneck in the vertical and hybrid algorithms.
With these limitations in mind, we propose Vertical-FLAFE as shown in Algorithm 2 and depicted in Figure 1. The basic idea for Vertical-FLAFE is that for features that are on separate clients and can’t be directly shared, we need to add Laplace noise to the vectors to be combined, homomorphically encrypt the vectors, combine the features in the encryped space, and finally send the combined features back to the clients to decrpyt. This combined feature can then be evaluated to see if it is a useful new feature.
@@
Once all of the engineered features are created, sent back to clients, and decrypted, the best engineered features are selected using any feature selection technique that abides by FL requirements. If significant computational resources are available, then a federated feature selection approach, such as the one we propose in Section 7.3 of the Appendix, can be used to very accurately select the best features. However, in the typical case of constrained resources, then a fully local feature selection that determines the best engineered features on each client separately, can be used. We use this approach in our experiments and provide details in Section 6.
It is important to note that most homomorphic encryption schemes require specifying the downstream function while encrypting the inputs, while in our approach we apply various functions (combinations of multiplications and additions) to the encrypted inputs. Fortunately, BFV and CKKS homomorphic encryption do not require the specification of the downstream function while encrypting and fit well within our approach [bfv1][bfv2][ckks].
We make some assumptions on the way that data is gathered across clients in the hybrid setting. While work such as [hyfdca] allows fairly flexible partitions of data as long as each client hold rectangular patches of the overall data, we enforce more strict rules on the way data is split. Specifically, we require that every sample be partitioned across clients in the same way. Essentially, the vertical divisions must run across the entire sample space without any jagged partitions. An example of a suitable partition is shown in Figure 2.
@@
Algorithm 4 demonstrates how feature combinations from different feature spaces are performed. In the vertical case, after the combination is formed, it is straightforward to randomly select one of the two clients that owned the original features to send the new feature to. However, in the hybrid case, the sample space is also partitioned, so it is possible that a combined feature $H^{1}_{i,j}$ is sent to the client holding original feature $i$ and later (in the different sample space) a combined feature $H^{2}_{i,j}$ is sent to the client holding original feature $j$. This would result in the overall data matrix becoming jagged and would cause errors in the subsequent feature selection and hybrid FL training. We handle this by keeping track of which group of clients that contain the same features we send this new feature to, and for subsequent combinations we assign the client to this same group to keep the feature space properly aligned.
1 Input: Clients $k=1,2,...,Q$, Local data matrix $X_{k}\in\mathbb{R}^{N_{k}\times M_{k}}$ and targets $y_{k}\in\mathbb{R}^{N_{k}}$ belonging to client $k$, CKKS Homomorphic encryption function $H(\cdot)$, $\text{Lap}(0|b)^{d}$ represents a vector of dimension $d$ with independent values each drawn from a Laplace distribution with scale $b$
@@ 5             $F_{out}=F_{out}\cup\text{bivariateFunction}(X_{k}[i],X_{k}[j])$
1 Input: Lists of encrypted feature vectors for each client denoted as $\mathcal{H}_{k}$, set of candidate bivariate transformations $\mathcal{B}_{nonlocal}$, define the set of client groups that share same sample space $\mathcal{G}_{s}$ (in vertical case this is a single group that contains all clients), define the set of client groups that share same feature space $\mathcal{G}_{f}$ (in vertical case each group will have a single client)
@@ 5             $\mathcal{H}=\mathcal{H}\cup\mathcal{H}_{k}$
8                   $H_{i,j}=\text{bivariateFunction}(H_{i},H_{j})$
@@ 14                         Send $H_{i,j}$ to $c$
16                         Select client that owns feature $i$ or feature $j$ at random with equal probability and send $H_{i,j}$ to selected client. Find feature client group $G_{c}$ that selected client belongs to
@@ 5 Experiments
We evaluate the performance of our algorithms in the horizontal, vertical, and hybrid FL settings. We test on the same eight datasets evaluated in [iife], which consist of four regression and four classification datasets. Details of these datasets are available in the Appendix Section 7.2. We use L2-regularized linear models for these experiments; logistic regression for classification problems and ridge regression for regression problems. We perform random search hyperparameter tuning before the AutoFE process and after the AutoFE process. For linear models, the only parameter tuned is the regularization constant. The hyperparameter tuning evaluations are performed in an FL fashion. Following the approach in [iife], our test metric is F1-micro for classification problems and 1 - (relative absolute error) for regression problems. For the linear models we use one-hot encoding of the categorical variables and standard scaling of the features when training and performing evaluations. We define $Q$ as the number of clients.
@@ 5.1.2 Verification of Federated Interaction Information
Computing interaction information values to predict the synergy of a feature pair with the target is a key step in IIFE. In [iife], the viability of interaction information was evaluated using a synthetic data experiment. The basic idea is to create a synthetic target as some function of two features. We would then expect the interaction information of the selected pair of features and the synthetic label to be high compared to other feature pairs. If we repeat this across many feature pairs in the dataset, in the optimal case, the true feature pair would be ranked 0 in terms of ranking by interaction information for every tested feature pair.
@@ 5.2.1 Algorithm Performance
For the Vertical FL AutoFE experiments, we create splits in the data by simply partitioning the feature space into equal segments. Due to each client having different features, the distributions of data across clients are non-IID by default. To protect client data from being reconstructed by the server or other clients, differential privacy is used when forming the feature combinations of features that belong on different clients as shown in Algorithm 2. We use additive Laplace noise which is $\epsilon$-differentially private [laplace]. We run experiments with no differential privacy and with differential privacy with $\epsilon=1$ and $\epsilon=0.1$, which are relatively low values indicating high privacy. We perform homomorphic encryption using the CKKS scheme [ckks] provided in the PyFHEL Python library [pyfhel]. We use HyFDCA as the FL training algorithm as vertical FL is a special case of hybrid FL and performs well in this setting [hyfdca]. Exact details of the experimental setup including differential privacy details are provided in Section 7.2 of the Appendix.

### v3

@@ 2 Introduction
Federated learning (FL) is a privacy-preserving machine learning technique that has gained traction in recent years. The main requirement in federated learning is that data belonging to client devices cannot be directly sent to the server to build a global model. Instead, the typical approach is for each client to build a local model with only their own local data, then the clients can send their model weights to the server which aggregates these weights in various ways. Many of the key challenges in federated learning revolve around minimizing the number of communications between the clients and server and maintaining good performance when the data across clients are non-IID Li et al., (2020).
There are three federated learning settings that correspond to how the data is gathered across clients. In horizontal federated learning, each client has a subset of samples and each of these samples has all of the features Li et al., (2020). In vertical federated learning, each client has the full set of samples but only a subset of the features Liu et al., (2022). Finally, in hybrid federated learning, each client only has a subset of samples and each of these samples only has a subset of features Overman et al., 2024a Zhang et al., (2020)Jang et al., (2024). Each of these settings requires different algorithmic approaches.
This paper focuses on automated feature engineering in the context of federated learning, where clients are not able to send data to the server, but we still build informative new features from the original raw features. We introduce algorithms for all three FL settings: horizontal, vertical, and hybrid. The horizontal algorithm we propose extends IIFE Overman et al., 2024b to the federated setting, which we term Fed-IIFE. The hybrid and vertical FL methods hinge on homomorphic encryption and differential privacy to safely form combinations of features that reside on different clients, along with a feature selection scheme to determine which features to keep. All of the algorithms are iterative in that they gradually construct more and more complicated engineered features each iteration.
@@ 3 Related Work
There is significant work existing on automated feature engineering. The main focus of most algorithms is on cleverly navigating the massive search space of possible combinations of features for high order engineered features. OpenFE uses a very fast boosting technique based on gradient boosted trees to quickly evaluate all possible feature combinations Zhang et al., (2023). AutoFeat enumerates all possible combinations up to a specified order, then uses a fast feature selection technique to select the best performing features Horn et al., (2020). IIFE tackles this problem by using interaction information to determine the best synergizing feature pairs and only exploring the best pairs of features, then iteratively builds higher order features Overman et al., 2024b . EAAFE uses a genetic algorithm approach to search the best engineered features Zhu et al., 2022a . DIFER uses a deep learning approach to find representations of the engineered feature strings and their predicted validation scores Zhu et al., 2022b . Furthermore, very recently there have been a focus on using Large Language Models to build engineered features in an informed fashion based on feature descriptions Hollmann et al., (2023).
@@
Federated learning also has extensive algorithmic work in addition to interesting theoretical results. Federated averaging is the original FL algorithm that is designed for the horizontal setting McMahan et al., (2017). Further improvements for heterogeneous data exist, such as SCAFFOLD Karimireddy et al., (2020) and RadFed Xue et al., (2022). The vertical FL setting has also had significant work in recent years Liu et al., (2022). The hybrid federated learning case where clients have subsets of both samples and features has far fewer algorithms, but has recently had more work exploring this case Overman et al., 2024a Zhang et al., (2020)Jang et al., (2024).
@@
Despite the great deal of work in federated learning and automated feature engineering, there is relatively limited work on automated feature engineering in the federated learning setting. All of the existing work at the time of this writing focuses on the vertical case, where each client has different features, and all clients have the same samples. There exists work in AutoML in a federated setting, however automated feature engineering is not one of the aspects discussed in this work Preuveneers, (2023). FLFE is an algorithm for vertical federated AutoFE that uses quantile sketching to encode features and uses deep neural networks to predict if a particular transformation of features is useful Fang et al., (2020). The neural networks are trained on a number of datasets, but these datasets are not shared, so it is impossible to reproduce their results. Since these datasets are actually used to build the feature engineering framework and not just used for evaluating the algorithm, it is necessary to have these datasets to properly evaluate the algorithm. Finally, homomorphic encryption is used when constructing the best candidate features for building the final model. However, this work shows only small gains compared to using only the raw features, does not compare to AutoFE methods in the centralized setting, and does not provide any source code making reproducibility very difficult. Furthermore, clients may be able to reconstruct the features of other clients from the decrypted engineered features; we address this in our work by utilizing differential privacy. We believe this makes the work very difficult to reproduce and compare. It is clear that there is a large research gap in all three settings of federated AutoFE: horizontal, vertical, and hybrid. Our work provides algorithms for all of these settings.
@@ 4 Algorithms
Fortunately, extending IIFE to the FL setting is relatively straightforward and results in good empirical performance compared to centralized IIFE, which is one of the top-performing AutoFE algorithms at the time of this writing (Overman et al., 2024b, ). We detail this algorithm, Fed-IIFE, in Section 4.1. The cases of vertical and hybrid FL are more complicated than horizontal FL and are discussed in Section 4.2.
@@ 4.1 Horizontal Setting
We extend IIFE to the FL setting by replacing some key aspects of IIFE with federated versions. We call this new FL version Fed-IIFE. Refer to the full algorithmic description of IIFE in Overman et al., 2024b . We replace computations of interaction information with federated interaction information described in Algorithm 1. The key change is that interaction information is computed separately on each client. The clients send these interaction information scores to the server; the server then averages the contributions from all of the clients. The averaging can either be the simple average or a weighted average where the weights are the proportion of samples belonging to the client. We use the weighted average in our experiments. The other key change from IIFE to Fed-IIFE is that all model evaluations are performed in a federated fashion. Candidate transformations are computed on each client separately and then evaluated across all clients using any horizontal FL algorithm. In our experiments, we use FedAvg (McMahan et al.,, 2017) for horizontal FL training.
@@ 4.2 Vertical and Hybrid Settings
Vertical AutoFE algorithms must significantly differ from centralized and horizontal FL settings due to the fact that each client holds different feature spaces. The key aspect of AutoFE is combining different features together in ways that improve downstream model performance. However, directly sharing features on different clients with each other or a central server directly violates the requirements of FL. Furthermore, many vertical and hybrid FL algorithms require homomorphic encryption to protect privacy leakages when sharing intermediate information (such as inner products) for computing gradients (Overman et al., 2024a, ). This is computationally expensive which severely limits the number of model training/evaluations that can be performed (AutoFE algorithms typically require a very large number of such trainings/evaluations). In fact, we show in Section 5.2.1 that homomorphic encryption is the main computational bottleneck in the vertical and hybrid algorithms.
With these limitations in mind, we propose Vertical-FLAFE as shown in Algorithm 2 and depicted in Figure 1. The basic idea for Vertical-FLAFE is that for features that are on separate clients and can’t be directly shared, homomorphically encrypt the vectors, combine the features in the encrypted space, add encrypted Laplace noise (the server can do this with a public key, but is not able to decrypt), and finally send the combined features back to the clients to decrypt. This combined feature can then be evaluated to see if it is a useful new feature. Full details of privacy considerations are provided in the Appendix.
@@
Once all of the engineered features are created, sent back to clients, and decrypted, the best engineered features are selected using any feature selection technique that abides by FL requirements. If significant computational resources are available, then a federated feature selection approach, such as the one we propose in Section 7.3 of the Appendix, can be used to very accurately select the best features. However, in the typical case of constrained resources, then a fully local feature selection that determines the best engineered features on each client separately can be used. We use this approach in our experiments and provide details in Section 6.
It is important to note that most homomorphic encryption schemes require specifying the downstream function while encrypting the inputs, while in our approach we apply various functions (combinations of multiplications and additions) to the encrypted inputs. Fortunately, BFV and CKKS homomorphic encryption do not require the specification of the downstream function while encrypting and fit well within our approach (Wibawa et al.,, 2022; Halevi et al.,, 2019; Cheon et al.,, 2017).
We make some assumptions on the way that data is gathered across clients in the hybrid setting. While hybrid training algorithms, such as that by Overman et al., 2024a , allow fairly flexible partitions of data as long as each client hold rectangular patches of the overall data, we enforce more strict rules on the way data is split. Specifically, we require that every sample be partitioned across clients in the same way. Essentially, the vertical divisions must run across the entire sample space without any jagged partitions. An example of a suitable partition is shown in Figure 2.
@@
Algorithm 4 demonstrates how feature combinations from different feature spaces are performed. In the vertical case, after the combination is formed, it is straightforward to randomly select one of the two clients that owned the original features to send the new feature. However, in the hybrid case, the sample space is also partitioned, so it is possible that a combined feature $H^{1}_{i,j}$ is sent to the client holding original feature $i$ and later (in the different sample space) a combined feature $H^{2}_{i,j}$ is sent to the client holding original feature $j$. This would result in the overall data matrix becoming jagged and would cause errors in the subsequent feature selection and hybrid FL training. We handle this by keeping track of which group of clients that contain the same features we send this new feature to, and for subsequent combinations we assign the client to this same group to keep the feature space properly aligned.
1 Input: Clients $k=1,2,...,Q$, Local data matrix $X_{k}\in\mathbb{R}^{N_{k}\times M_{k}}$ and targets $y_{k}\in\mathbb{R}^{N_{k}}$ belonging to client $k$, CKKS Homomorphic encryption function $H(\cdot)$
@@ 5             $F_{out}=F_{out}\cup\text{bivariateFunction}(X_{k}[i],X_{k}[j])$
1 Input: Lists of encrypted feature vectors for each client denoted as $\mathcal{H}_{k}$, CKKS Homomorphic encryption function $H(\cdot)$, $\text{Lap}(0|b)^{d}$ represents a vector of dimension $d$ with independent values each drawn from a Laplace distribution with scale $b$, set of candidate bivariate transformations $\mathcal{B}_{nonlocal}$, define the set of client groups that share same sample space $\mathcal{G}_{s}$ (in vertical case this is a single group that contains all clients), define the set of client groups that share same feature space $\mathcal{G}_{f}$ (in vertical case each group will have a single client)
@@ 5             $\mathcal{H}=\mathcal{H}\cup\mathcal{H}_{k}$
                   $H_{i,j}=\text{bivariateFunction}(H_{i},H_{j})+H(\text{Lap}(0|b)^{N_{k}})$ // inject encrypted Laplace noise
@@ 13                         Send $H_{i,j}$ to $c$
15                         Select client that owns feature $i$ or feature $j$ at random with equal probability and send $H_{i,j}$ to selected client. Find feature client group $G_{c}\in\mathcal{G}_{f}$ that selected client belongs to
@@ 5 Experiments
We evaluate the performance of our algorithms in the horizontal, vertical, and hybrid FL settings. We test on the same eight datasets evaluated in Overman et al., 2024b , which consist of four regression and four classification datasets. Details of these datasets are available in the Appendix Section 7.2. We use L2-regularized linear models for these experiments; logistic regression for classification problems and ridge regression for regression problems. We perform random search hyperparameter tuning before the AutoFE process and after the AutoFE process. For linear models, the only parameter tuned is the regularization constant. The hyperparameter tuning evaluations are performed in an FL fashion. Following the approach in Overman et al., 2024b , our test metric is F1-micro for classification problems and 1 - (relative absolute error) for regression problems. For the linear models we use one-hot encoding of the categorical variables and standard scaling of the features when training and performing evaluations. We define $Q$ as the number of clients.
@@ 5.1.2 Verification of Federated Interaction Information
Computing interaction information values to predict the synergy of a feature pair with the target is a key step in IIFE. In Overman et al., 2024b , the viability of interaction information was evaluated using a synthetic data experiment. The basic idea is to create a synthetic target as some function of two features. We would then expect the interaction information of the selected pair of features and the synthetic label to be high compared to other feature pairs. If we repeat this across many feature pairs in the dataset, in the optimal case, the true feature pair would be ranked 0 in terms of ranking by interaction information for every tested feature pair.
@@ 5.2.1 Algorithm Performance
For the Vertical FL AutoFE experiments, we create splits in the data by simply partitioning the feature space into equal segments. Due to each client having different features, the distributions of data across clients are non-IID by default. To protect client data from being reconstructed by the server or other clients, differential privacy is used when forming the feature combinations of features that belong on different clients as shown in Algorithm 2. We use additive Laplace noise which is $\epsilon$-differentially private Dwork et al., (2006). We run experiments with no differential privacy and with differential privacy with $\epsilon=1$ and $\epsilon=0.1$, which are relatively low values indicating high privacy. We perform homomorphic encryption using the CKKS scheme Cheon et al., (2017) provided in the PyFHEL Python library Ibarrondo and Viand, (2021). We use HyFDCA as the FL training algorithm as vertical FL is a special case of hybrid FL and performs well in this setting Overman et al., 2024a . Exact details of the experimental setup including differential privacy details are provided in Section 7.2 of the Appendix.